OpenStudy (anonymous):

2^(x+2) + 2^x = 320 .. what is the answer for x?

OpenStudy (anonymous):

\[2^{x+2}=2^22^x=4\cdot2^x\]

OpenStudy (raden):

2^(x+2) + 2^x = 320 2^2 * 2^x + 2^x = 320 4 * 2^x + 2^x = 320 5 * 2^x = 320 divide by 5 to both sides, giving us 2^x = 64 2^x = 2^6 so, x = .....

OpenStudy (anonymous):

whats your answer raden.. and i don't think thats the ritght way of doing it

OpenStudy (anonymous):

@RadEn didn't give you the answer, but it should be clear what it is. All his work his right.

OpenStudy (raden):

in the last step, cancel the base 2 from both sides, take the rest as ur answr

OpenStudy (anonymous):

ohh okay makes sense.. x is 6!

OpenStudy (raden):

haha, good finishing :)

OpenStudy (anonymous):

thank you so much :)! i have another question could i just ask it here?

OpenStudy (anonymous):

log(base 2) (3x+2) + log(base2) (x-1) = 3